# Robust Containment Queries over Collections of Rational Parametric Curves via Generalized Winding Numbers

### v1

@@ Abstract.
Point containment queries for regions bound by watertight geometric surfaces, i.e. closed and without self-intersections, can be evaluated straightforwardly with a number of well-studied algorithms. However, when such assumptions on domain geometry are not met, these methods are theoretically unfounded at best and practically unusable at worst. More robust classification schemes utilize generalized winding numbers, a mathematical construction that is indifferent to imperfections in the often human-defined geometric model. We extend this methodology to more general curved shapes, defining a robust containment query for regions whose boundary elements are defined by a collection of rational parametric curves. In doing so, we devise an algorithm that is stable and accurate at arbitrary points in space, circumventing the typical difficulties for queries that are arbitrarily close or coincident with the model. This is done by reducing the generalized winding number problem to an integer winding number problem, which is solved by approximating each curve with a polyline that provably has the same winding number at the point of interest. We demonstrate the improvements in computational complexity granted by this method over conventional techniques, as well as the robustness induced by its application.
@@ 1. Introduction
As a motivating example, consider multiphysics simulations involving the interactions of several geometrically complicated shapes within a computational domain on which the interaction of governing PDEs are modeled. Many traditional approaches construct a conforming volumetric mesh directly from a boundary description (Thompson et al., 1985), but this is itself a complex task, as the generated mesh must have high enough quality to accurately represent both the geometry present and the numerical solution of the physical system (Ingram et al., 2003).
Figure 2. Representations of a geometric object: A boundary representation (left), a volumetric representation generated by the boundary (center) and a boundary representation embedded in a computational grid (right).
Generation of a direct volumetric representation poses additional difficulties in the presence of a curved B-Rep. In such cases, most volumetric meshes generated for this purpose are still composed of linear elements, which is a known, yet often accepted, source of error in such problems. We refer to Sevilla et al. (2008) for a more comprehensive overview of the complications that arise when a piecewise linear boundary is used as an approximate of a curve, but it necessarily leads to loss of curvature and other geometric features unless excessive and expensive refinement of the mesh is used. In the context of finite element analysis, this has led to the development of techniques that work directly on the boundary as described by the underlying CAD geometry (Hafner et al., 2019). As widespread categories of examples, NURBS-enhanced finite element analysis allows the boundary of a computational cell to share a curve with the geometric model (Sevilla et al., 2008, 2011; Sevilla and Huerta, 2018), and methods in isogeometric analysis operate on the boundary model without reference to the spatial discretization itself (Hughes et al., 2005; Marussig and Hughes, 2017). Particularly relevant to the forthcoming discussion are immersed boundary (Peskin, 2002) and immersogeometric (Kamensky et al., 2015) methods, which solve problems in fluid-structure interaction by placing the complicated CAD model on a simpler computational grid along which the numerical solution is computed.
To effectively utilize many of these techniques, it is necessary to determine whether points at arbitrary locations within the domain are contained by the boundary representation, as this affects local physical and material properties. For example, multimaterial Arbitrary Lagrangian-Eulerian (ALE) methods require additional preprocessing to initialize the volume fraction of each material in a computational cell (Hirt et al., 1974; Barlow et al., 2016). This, in turn, necessitates classifying each point in a quadrature rule along the background grid as inside or outside the volume bounded by the corresponding shape (Weiss et al., 2016). Although the positions of these query points might be structured from the perspective of the background grid, they are arbitrary relative to the shape geometry itself. It is critical that this identification takes into account the full curvature of the bounding geometry, as piecewise linear approximations can cause quadrature points near material boundaries to be misclassified, unpredictably compounding the error in the calculation. It is in this context that we focus our attention on the containment query, a geometric predicate that returns whether an arbitrary spatial location is contained within an arbitrary shape.
@@
Owing to advances in the field of computer graphics, the rapid evaluation of containment queries is quite well understood for certain classes of shapes, such as polygons and Bézier curves (Nishita et al., 1990; Carvalho and Cavalcanti, 1995; Patrikalakis and Maekawa, 2010). However, such methods assume the boundary is watertight, having no gaps between connected components. In reality, the vast majority of B-Reps are created by hand within some form of CAD software, leading to so-called “messy” CAD geometry. A straightforward source of these problems is human error during construction, but more subtle issues can be induced through varying tolerances within and between different software tools. This results in boundary models with human-imperceptible, but numerically significant gaps and overlaps between individual components of the boundary (Marussig and Hughes, 2017) (see Figure 3).
Even in situations where visibility of the model is the primary concern, such as in computer graphics and animation, care must be taken to ensure that these geometric errors are handled properly. For example, the dynamic planar map heuristic used by Adobe Illustrator’s Live Paint Bucket tool allows users to set fills for closed regions bounded by intersecting collections of Bézier curves and utilizes a user-provided threshold to assist with connecting gaps between curves (Asente et al., 2007). These geometric errors are even less acceptable when the model is used in the context of scientific simulation. The mathematical underpinnings of containment within non-watertight geometries is dubious as such objects have no interior or exterior in a strict sense. This leads to a tenuous theoretical foundation for established methods that assume geometric continuity, but more practically, it often results in unexpected and fatal errors in simulation software that wastes computational resources and development time.
@@
Such geometric errors can be resolved with techniques in surface repair (Mezentsev and Woehler, 1999; Park et al., 2021; Bischoff et al., 2005), which alter the B-Rep itself to an approximated surface that lacks these imperfections. However, doing so by hand can be impractical for the typical number of imperfections in such a model, and automated procedures during geometry pre-processing can inadvertently remove important model features. Instead, we require a containment query that is robust to such artifacts, such that an appropriate inside/outside classification is always returned, even in the presence of these, possibly imperceptible, surface errors.
In this work, we present three principal contributions.
  • We extend the theoretical framework of generalized winding numbers (Jacobson et al., 2013) to the context of an unstructured collection of 2D curves, from which one can derive a robust containment query that indicates whether a point should be considered interior to non-watertight geometry.
@@
  • We provide a novel algorithm for evaluating integer winding numbers with respect to curved geometry, which is the principal component of our technique for evaluating generalized winding numbers on such shapes. We demonstrate this algorithm to be considerably more performant than current state-of-the-art techniques.
Altogether, these contributions allow for an implicit volumetric description of an object implied by messy 2D CAD geometry, composed of collections of parametric curves (see Figure 1). Our algorithm has been implemented in Axom, a BSD-licensed open-source library of Computer Science infrastructure for HPC applications (Capps et al., 2024).
@@ 2. Background and Related Work
Ray Casting algorithms determine containment by extending a ray from the specified query point out to infinity, and counting the number of times the ray intersects the domain boundary to produce a crossing number (Shimrat, 1962). As a point moves along this ray, it will alternate between interior and exterior, and therefore an odd (even) crossing number indicates the point is interior (exterior). Computing intersections between the given ray and the linear edges of a polygon can be done very efficiently, with many graphics processing units (GPUs) having dedicated ray-tracing cores that can perform the operation massively in parallel (Zellmann et al., 2022). However, these methods are sensitive to the specific ray that is extended, needing to account for special cases when the ray intersects a polygon vertex or entire edge, which can be unnecessarily more costly (Edelsbrunner and Mücke, 1990; Sommariva and Vianello, 2022).
Winding number algorithms are a class of algorithms that are more indifferent to these issues, and, as such, are the type from which we derive our proposed method. In a winding number algorithm, we instead count the number of revolutions around the query point made by a particle traveling on the domain boundary (Hormann and Agathos, 2001). This value can be considered a generalization of the crossing number, and can be used to determine containment with the same even-odd rule. By virtue of not being dependent on any particular extended ray, these algorithms implicitly handle edge-cases that are problematic for ray casting. This means even in the context of simple, watertight geometry, such algorithms can be considered a more robust approach to containment queries.
@@
Containment of arbitrary points in more general curved shapes is a much more difficult problem, and has been studied extensively in the context of 2D vector graphics. As an example, we refer to the process of linearization, in which a curve is approximated by connected linear segments, an approach that is fairly common in both modern graphics engines and physics codes. In the former, information about the targeted display can be leveraged to ensure that the approximating polyline is constructed with sufficiently high resolution so as to be indistinguishable to a viewer (Kilgard, 2020). Combined with the fact that the polyline must be rendered with some predetermined thickness, it is therefore theoretically possible (although rarely enforced) for containment queries on linearized watertight geometries on a rendered display to have no observable misclassifications.
@@
We similarly desire an algorithm that can reach this level of geometric fidelity for a very general class of curves. For example, cubic Bézier curves are the highest order required by postscript and SVG formats, as the level of detail they afford is typically sufficient for CAD applications. However, applications in computational multiphysics can necessitate an algorithm that works for rational Bézier curves of arbitrary order. This is because the NURBS shapes that are used by more sophisticated modeling software can always be decomposed into a collection of (in principle arbitrary order) rational Bézier primitives through the process of Bézier extraction (Piegl and Tiller, 1996; Thomas et al., 2015). However as we will see, the proposed algorithm is ultimately indifferent to the order and/or rationality of the curve in question.
@@
These issues are worsened by geometric errors present in the shape. As an example, if the arbitrary ray is by chance extended through a very small gap in the model, then the point must be classified as “exterior” despite this being unintuitive and contrary to the intentions of the mesh designer. Non-manifold edges in a shape are similarly problematic for ray casting methods. These issues can be somewhat mitigated by extending several rays and taking a consensus, but such approaches impose an additional computational burden while still producing noisy and potentially inconsistent classifications for nearby points (Nooruddin and Turk, 2003). We now show that, in addition to being robust to the specific edge cases introduced by ray casting, winding number algorithms can be extended such that they are also robust to more general issues introduced by non-watertight, non-manifold, curved geometry.
@@ 3. Generalized Winding Numbers
The generalized winding number is an extension of the standard integer-valued winding number to (potentially) non-watertight regions. While integer winding numbers partition the enclosed regions of a domain, the generalized winding number generates a harmonic scalar field that smoothly degrades in the presence of discontinuities and self-intersections allowing for robust containment queries in the presence of messy geometry. As an example, a simple rule for handling the fractional values that result from imperfections in the bounding geometry is to round them to their nearest integer.
@@
Such generalizations of winding numbers have recently shown great utility in geometry processing applications. In the context of both (linear) triangle meshes (Jacobson et al., 2013) and point clouds (Barill et al., 2018), the resulting inside/outside classification can be used to generate volumetric triangle/tetrahedral meshes out of completely unstructured geometric data. Additionally, the smooth degradation of the winding number field naturally introduces desirable locality properties, which in turn lead to more robust and performant algorithms in, for example, exact mesh booleans (Trettner et al., 2022). More recently, these particular winding number algorithms have been used in the context of immersogeometric analysis to simulate fluid flow around geometric objects defined by unstructured point clouds (Balu et al., 2023). It is speculated in Gunderman (2021) that one could extend the calculation of generalized winding numbers to curved geometry by utilizing their own novel quadrature schemes for such objects (Gunderman et al., 2020), but preliminary tests of this approach is still limited by the accuracy of numerical integration. Despite this recent surge in use, methods that compute exact generalized winding numbers for more general curved geometric objects are, to our knowledge, currently absent from the literature.
@@
Nevertheless, the problem of computing generalized winding numbers for curved shapes has close analogues in the fields of both computer graphics and boundary integral equations. In the former, a connection can be drawn to diffusion curves, a primitive in vector graphics that also produces a scalar field defined by the solution of a differential equation with respect to boundary curves (albeit with different data prescribed on the curve) (Jacobson et al., 2013). These applications have the advantage of being solved on a grid generated by a rasterized domain, meaning that a global solution can be obtained efficiently using a geometric multi-grid method (Orzan et al., 2008). Sawhney and Crane (2020) recently introduced a grid-free approach in which random walks and projections onto curved geometry are used to solve certain classes of PDEs, including that which implicitly defines the generalized winding number, focusing on improving speed and locality at the cost of some degree of accuracy and consistency. Within the context of boundary integral equations, calculating winding numbers is a special case of the more general problem of integrating double layer potentials with a constant density function. However, such algorithms are typically applied to closed domains with simpler (although still curved) types of geometry, precluding general use on arbitrary, messy CAD models (Klinteberg and Barnett, 2019). Furthermore, our proposed algorithm relies exclusively on geometric principles, eliminating the need to use costly and potentially inaccurate quadrature schemes.
@@
We begin by reviewing the formal description of a generalized winding number in $\mathbb{R}^{2}$. Given an oriented curve $\Gamma\subset\mathbb{R}^{2}$ and query point ${q}\in\mathbb{R}^{2}\setminus\Gamma$, the winding number $w$ describes the (potentially incomplete) number of times the curve travels counterclockwise around the point. In the case where $\Gamma$ is a closed curve, the scalar field $w_{\Gamma}({q})$ is integer-valued, and induces a partition of $\mathbb{R}^{2}$ that can be interpreted as containment in the region bounded by $\Gamma$. In the following discussion, it is notationally convenient to assume that the query point $q$ is located at the origin, and that the curve $\Gamma$ is similarly translated by $q$, as the scalar field is invariant to such global translations.
@@
When $\Gamma$ is piecewise linear, i.e. $\Gamma=\cup_{i}L_{i}$, the winding number at $q$ can be computed as the sum of signed angles $\theta_{i}$ subtended by each linear component $L_{i}$, given by
@@
It is from this property of the winding number that we derive the robustness of our containment queries. While the collection of Bézier curves may be messy in the sense that pairs of endpoints may not perfectly align as is the intention of the designer, generalized winding numbers on each individual curve are well-defined. The sum of these values at a given point is the generalized winding number of the entire shape, which can be used to produce accurate and intuitive containment queries (see Figure 6). Most commonly, this is done by rounding this value to the nearest integer, and applying to it a conventional even-odd or non-zero rule to make the classification. Importantly, each of these conventions treat points with a zero winding number as exterior.
@@
Direct evaluation of this integral through standard techniques is notoriously difficult. In particular, when $\Gamma$ is close to the query point (or rather the origin under this translated notation) then the near-singular behavior of the integrand causes Gaussian quadrature to become unstable, as seen in Figure 7. Although some error is to be expected for this type of numerical integration, this error rapidly becomes unacceptable as the query point approaches the curve. Perhaps more concerningly, there is no immediate way to verify that the value returned by the method is accurate. We can see how this would affect a containment query on a closed shape in Figure 7(middle, bottom). In this example, the only correct values for the rounded winding number are 0, 1, and 2. However, even the most accurate quadrature scheme shown produces values in the vicinity of the quadrature nodes that are completely meaningless.

### v2

@@ Abstract.
Point containment queries for regions bound by watertight geometric surfaces, i.e., closed and without self-intersections, can be evaluated straightforwardly with a number of well-studied algorithms. When this assumption on domain geometry is not met, such methods are either unusable, or prone to misclassifications that can lead to cascading errors in downstream applications. More robust point classification schemes based on generalized winding numbers have been proposed, as they are indifferent to these imperfections. However, existing algorithms are limited to point clouds and collections of linear elements. We extend this methodology to encompass more general curved shapes with an algorithm that evaluates the winding number scalar field over unstructured collections of rational parametric curves. In particular, we evaluate the winding number for each curve independently, making the derived containment query robust to how the curves are arranged. We ensure geometric fidelity in our queries by treating each curve as equivalent to an adaptively constructed polyline that provably has the same generalized winding number at the point of interest. Our algorithm is numerically stable for points that are arbitrarily close to the model, and explicitly treats points that are coincident with curves. We demonstrate the improvements in computational performance granted by this method over conventional techniques as well as the robustness induced by its application.
@@
[…]
Figure 1: A comparison between containment queries using traditional ray casting and generalized winding numbers is shown for a simple vector graphic. The ray casting method fails to correctly classify points near gaps, overlaps, or non-maniforld edges, while the field of generalized winding numbers is smooth around these artifacts. Three methods of computing generalized winding numbers are shown on a zoomed-in portion of the figure, including an unstable quadrature, an inaccurate linear approximation, and our proposed adaptive algorithm.
@@ 1. Introduction
Such a task motivates development of the containment query, a geometric predicate that returns whether an arbitrary spatial location is contained within an arbitrary shape. The rapid evaluation of containment queries is quite well understood for many classes of shapes, such as polygons and regions bounded by Bézier curves (Nishita et al., 1990; Carvalho and Cavalcanti, 1995; Patrikalakis and Maekawa, 2010). However, such methods assume the boundary is watertight, having no gaps between connected components. In reality, the vast majority of B-Reps are created by hand within some form of CAD software, leading to so-called “messy” CAD geometry.
@@
[…]
Figure 2: We zoom in on a vertex of a simple 2D shape, which reveals that the model is unexpectedly non-watertight.
A straightforward source of these problems is human error during design or construction, but more subtle issues can be introduced through varying tolerances within and between different software tools. This results in boundary models with human-imperceptible, but numerically significant gaps and overlaps between individual components of the boundary (Marussig and Hughes, 2017) (see Figure 2). Even in situations prioritizing visibility over strict geometric fidelity, such as in computer graphics and animation, care must be taken to ensure that these errors are properly accounted for. For example, the dynamic planar map heuristic used by Adobe Illustrator’s Live Paint Bucket tool allows users to set fills for closed regions bounded by intersecting collections of Bézier curves and exposes a global threshold parameter to connect gaps between curves in the collection (Asente et al., 2007).
As an additional class of motivating examples, we consider applications in multiphysics simulation, for which containment classification errors can lead to cascading problems and incorrect results. The mathematical underpinnings of containment within non-watertight shapes is dubious, which leads to a tenuous theoretical foundation for established methods that assume geometric continuity. This is problematic, as errors are frequently inconsistent within local regions, with even arbitrarily close points receiving different classifications. More practically, these errors can result in unexpected and fatal errors in simulation software, wasting computational resources and developer time (Sederberg et al., 2008).
At the same time, physical applications often prevent approximating curved geometry with piecewise linear segments. Although this would allow for treatment of the messy—but now linear—geometry with established methods, such an approximation can produce sizable errors in the relevant numerical method (Sevilla et al., 2008). As a concrete example, multimaterial Arbitrary Lagrangian-Eulerian (ALE) methods require additional preprocessing to initialize the volume fraction of each material in a computational cell (Hirt et al., 1974; Barlow et al., 2016). This, in turn, necessitates classifying each point in a quadrature rule along the background grid as inside or outside the volume bounded by the corresponding shape (Weiss et al., 2016). It is critical that this identification takes into account the full curvature of the bounding geometry, as piecewise linear approximations can cause quadrature points near material boundaries to be misclassified, unpredictably compounding the error in the calculation.
Such geometric errors can be resolved with techniques in surface repair (Mezentsev and Woehler, 1999; Park et al., 2021; Bischoff et al., 2005), which alter the B-Rep itself to an approximated surface that lacks these imperfections. However, doing so by hand can be impractical for the typical number of imperfections in such a model, and automated procedures during geometry pre-processing can inadvertently remove important model features. Instead, we require a containment query that is robust to such artifacts, in the sense that an appropriate inside/outside classification is always returned, independent of size and frequency of these boundary errors.
To address these concerns, we present in this work three principal contributions:
  • We extend the theoretical framework of generalized winding numbers (Jacobson et al., 2013) to the context of an unstructured collection of 2D rational parametric curves, from which one can derive a robust containment query that indicates whether an arbitrary point should be considered interior to non-watertight geometry.
@@
  • We provide a novel algorithm for evaluating integer winding numbers with respect to curved geometry, which is the principal component of our technique for evaluating generalized winding numbers on such shapes. We demonstrate that this calculation is robust to non-watertight geometry and considerably more performant than state-of-the-art techniques.
Altogether, these contributions allow for robust containment queries of objects implied by messy 2D CAD geometry, composed of collections of parametric curves (see Figure 1). Our algorithm has been implemented in Axom, a BSD-licensed open-source library of Computer Science infrastructure for HPC applications (Capps et al., 2024).
@@ 2. Background and Related Work
Ray Casting algorithms determine containment by extending a ray from the specified query point out to infinity, and counting the number of times the ray intersects the domain boundary to produce a crossing number (Shimrat, 1962). As a point moves along this ray, it will alternate between interior and exterior, and therefore an odd (even) crossing number indicates the point is interior (exterior). Computing intersections between the given ray and the linear edges of a polygon can be done very efficiently, with many graphics processing units (GPUs) having dedicated ray-tracing cores that can perform the operation massively in parallel (Zellmann et al., 2022). However, these methods are sensitive to the specific ray that is extended, needing to account for special cases when the ray intersects a polygon vertex or entire edge, which can be unnecessarily more costly (Edelsbrunner and Mücke, 1990; Sommariva and Vianello, 2022). Winding number algorithms are a class of algorithms that are more indifferent to these issues, and, as such, are the type from which we derive our proposed method. In a winding number algorithm, we instead count the number of revolutions around the query point made by a particle traveling on the domain boundary (Hormann and Agathos, 2001). This value can be considered a generalization of the crossing number, and can be used to determine containment with the same even-odd rule. By virtue of not being dependent on any particular extended ray, these algorithms implicitly handle edge-cases that are problematic for ray casting. This means even in the context of simple, watertight geometry, such algorithms can be considered a more robust approach to containment queries.
@@
[…]
Figure 3: Fully described in the caption.
Containment of arbitrary points in more general curved shapes is a much more difficult problem, and has been studied extensively in the context of 2D vector graphics. As an example, we refer to the process of linearization, in which a curve is approximated by connected linear segments, an approach that is fairly common in both modern graphics engines and physics codes. In the former, information about the targeted display can be leveraged to ensure that the approximating polyline is constructed with sufficiently high resolution such that it is indistinguishable to a viewer (Kilgard, 2020). Combined with the fact that the polyline must be rendered with some predetermined thickness, it is therefore theoretically possible (although rarely enforced) for containment queries on linearized watertight geometries on a rendered display to have no observable misclassifications.
@@
We similarly desire an algorithm that can reach this level of geometric fidelity for a very general class of curves. For example, cubic Bézier curves are the highest order required by postscript and SVG formats, as the level of detail they afford is typically sufficient for design applications. However, applications in computational multiphysics can necessitate an algorithm that works for rational Bézier curves of arbitrary order. This is because the NURBS shapes that are used by more sophisticated modeling software can always be decomposed into a collection of (in principle arbitrary order) rational Bézier primitives through the process of Bézier extraction (Piegl and Tiller, 1996; Thomas et al., 2015). However as we will see, the proposed algorithm is ultimately indifferent to the order and/or rationality of the curve in question.
@@
These issues are worsened by geometric errors present in the shape. As an example, if the arbitrary ray is by chance extended through a very small gap in the model, then the point must be classified as “exterior” despite this being unintuitive and contrary to the intentions of the mesh designer. Non-manifold edges in a shape are similarly problematic for ray casting methods. These issues can be somewhat mitigated by extending several rays and taking a consensus, but such approaches impose an additional computational burden while still producing noisy and potentially inconsistent classifications for nearby points (Nooruddin and Turk, 2003). We now show that, in addition to being indifferent to the specific edge cases introduced by ray casting, winding number algorithms can be extended such that they are also robust to more general issues introduced by non-watertight, non-manifold, curved geometry.
@@ 3. Generalized Winding Numbers
The generalized winding number is an extension of the standard integer-valued winding number to (potentially) non-watertight regions. While integer winding numbers partition the enclosed regions of a domain, the generalized winding number generates a harmonic scalar field that smoothly degrades in the presence of discontinuities and self-intersections. This allows for the calculation of robust containment queries in the presence of messy geometry. As an example, a simple rule for handling the fractional values that result from imperfections in the bounding geometry is to round them to their nearest integer.
@@
Such generalizations of winding numbers have recently shown great utility in geometry processing applications. In the context of both (linear) triangle meshes (Jacobson et al., 2013) and point clouds (Barill et al., 2018), the resulting inside/outside classification can be used to generate volumetric triangle/tetrahedral meshes out of completely unstructured geometric data. Additionally, the smooth degradation of the winding number field naturally introduces desirable locality properties, which in turn lead to more robust and performant algorithms in, for example, exact mesh booleans (Trettner et al., 2022). More recently, these particular winding number algorithms have been used in the context of immersogeometric analysis to simulate fluid flow around geometric objects defined by unstructured point clouds (Balu et al., 2023). It is speculated in Gunderman (2021) that one could extend the calculation of generalized winding numbers by utilizing their own novel quadrature schemes for such objects (Gunderman et al., 2020), but preliminary tests of this approach are limited by the accuracy of numerical integration. Despite this recent surge in use, methods that compute exact generalized winding numbers for more general curved geometric objects are, to our knowledge, currently absent from the literature.
@@
Nevertheless, the problem of computing generalized winding numbers for curved shapes has close analogues in the fields of both computer graphics and boundary integral equations. In the former, a connection can be drawn to diffusion curves, a primitive in vector graphics that also produces a scalar field defined by the solution of a differential equation with respect to boundary curves (albeit with different data prescribed on the curve) (Jacobson et al., 2013). These applications have the advantage of being solved on a grid generated by a rasterized domain, meaning that a global solution can be obtained efficiently using a geometric multi-grid method (Orzan et al., 2008). Sawhney and Crane (2020) recently introduced a grid-free approach in which random walks and projections onto curved geometry are used to solve certain classes of PDEs, including that which implicitly defines the generalized winding number, focusing on improving speed and locality at the cost of some degree of accuracy and consistency. Within the context of boundary integral equations, calculating winding numbers is a special case of the more general problem of integrating double layer potentials with a constant density function. However, such algorithms are typically applied to closed domains with simpler (although still curved) types of geometry, precluding general use on arbitrary, messy CAD models (Klinteberg and Barnett, 2020). Furthermore, our proposed algorithm relies exclusively on geometric principles, eliminating the need to use costly and potentially inaccurate quadrature schemes.
@@
We begin by reviewing the formal description of a generalized winding number in $\mathbb{R}^{2}$. Given an oriented curve $\Gamma\subset\mathbb{R}^{2}$ and query point ${q}\in\mathbb{R}^{2}\setminus\Gamma$, the winding number $w$ describes the (potentially incomplete) number of times the curve travels counterclockwise around the point. In the case where $\Gamma$ is a closed curve, the scalar field $w_{\Gamma}({q})$ is integer-valued, and induces a partition of $\mathbb{R}^{2}$ that can be interpreted as containment in the region bounded by $\Gamma$. In the following discussion, it is notationally convenient to assume that the query point is located at the origin, and that the curve $\Gamma$ has been appropriately translated by $q$. The scalar field is invariant to such global translations, and so the winding number of $\Gamma$ at $q$, $w_{\Gamma}(q)$, is identical to the winding number at the origin after translating the curve, $w_{\Gamma-q}(0)$.
@@
When the curve is piecewise linear with $\Gamma=\cup_{i}L_{i}$, the winding number at $q$ can be computed as the sum of signed angles $\theta_{i}$ subtended by each linear component $L_{i}$, given by
@@
[…]
Figure 4: The winding number field for a closed polygon is shown to be the sum of the angles subtended by each edge at the point, which is always an integer. The field is then shown for a single edge of the polygon, which is still equal to the angle subtended by the edge at each point in the domain, but no longer an integer anywhere.
@@
It is from this property of the generalized winding number that we derive both the robustness of our winding number calculation, and the robustness of the derived containment query. While the collection of Bézier curves may be messy in the sense that pairs of endpoints may not perfectly align, generalized winding numbers on each individual curve are well-defined and can be computed stably. At the same time, the sum of these values at a given point is exactly the generalized winding number of the entire shape, which can be used to produce intuitive and robust containment queries for non-watertight shapes (see Figure 5). Most commonly, this is done by rounding this value to the nearest integer, and applying to it a conventional even-odd or non-zero rule to make the classification. Importantly, each of these conventions treat points with a zero winding number as exterior.
@@
[…]
Figure 5: The generalized winding number field is shown for a simple heart shape, which is equal to the integer 1 for all interior points and 0 for all exterior points. The field is then shown for the four curves that compose the shape, which are each computed independently of the others. Finally, the field is shown for the same curves reassembled into a shape that is similar to the original heart, but is not watertight, for which the field is approximately 1 on the interior and approximately 0 on the exterior.
@@
Direct evaluation of this integral through standard techniques is notoriously difficult. In particular, when $\Gamma$ is close to the query point (or rather the origin under this translated notation) then the near-singular behavior of the integrand causes Gaussian quadrature to become unstable, as seen in Figure 6. Although some error is to be expected for this type of numerical integration, this error rapidly becomes unacceptable as the query point approaches the curve. Perhaps more concerningly, there is no immediate way to verify that the value returned by the method is accurate. We can see how this would affect a containment query on a closed shape in Figure 6. In this example, the only correct values for the rounded winding number are 0, 1, and 2. However, even the most accurate quadrature scheme shown produces values in the vicinity of the quadrature nodes that are completely meaningless.
@@
[…]
Figure 6: Large numerical error appears when the winding number is computed via quadrature for points near the curve, which is shown on a simple curve. On a simple non-manifold shape, the consequences of this inaccuracy on containment classifications are shown, with even a high-order quadrature rule resulting in many misclassifications.